What is -8-(3x+6)=4-x?
Answer by Timnewman(323)

Hi dear,
The term -8-(3x+6)=4-x,is a linear equation in one variable.
The variable is x.If you wish to find the value of x,you can solve it as follows:
-8-3x-6=4-x
-14-3x=4-x
-3x+x=4-14
-2x=-10
x=5
Hence,the value of x is 5
Hope this helps?